OpenStudy (anonymous):

-2/3√63x^3 Simplify.

OpenStudy (anonymous):

Is this you question: \[-\frac{2}{\sqrt{63x^3}}\] or \[-\frac{2}{\sqrt{63}x^3}\]

OpenStudy (anonymous):

its -2/3 <-- fraction then √63x^3

OpenStudy (anonymous):

Can you do the whole thing using the equation tool in post because I still don't understand. And are the instructions to rationalize the numerator?

OpenStudy (anonymous):